You have just graduated from the Information Communication University (ICU) of

Lusaka in Zambia and are leaving on a whirlwind tour to see some friends. You wish to

spend USD 1,000 each in Germany, New Zealand, and Great Britain (USD 3,000 in total).

Your bank offers you the following bid-ask quotes:

USD/EUR 1.304-1.305, USD/NZD 0.67­0.69, and USD/GBP 1.90-1.95.


(a)      If you accept these quotes, how many EUR, NZD, and GBP do you have at departure?                                                                                                 

(b)      If you return with EUR 300, NZD 1,000, and GBP 75, and the exchange rates are unchanged, how many USD do you have?                                          

(c)      Suppose that instead of selling your remaining EUR 300 once you return home, you want to sell them in Great Britain. At the train station, you are offered GBP/EUR 0.66-0.68, while a bank three blocks from the station offers GBP/EUR 0.665-0.675.                                                                                                       

 i) At what rate are you willing to sell your EUR300?                               

          ii) How many GBP will you receive? 


1
Expert's answer
2021-11-17T10:03:01-0500

a.

The currency can be converted into another currency through banks or forex dealers. The rate they quote is called a bid-ask quote.

The bid price is the one at which the bank is willing to buy the currency.

The ask price is the price at which bank is willing to sell the currency.

USD to EUR

1000 / 1.305 = 766.2835


USD to NZD

1000 / 0.69 = 1449.2754


USD to GBP

1000 / 1.95 = 512.8205


b.

 I leave with some currency and that I want to convert into the USD.

This will be converted at the bid rate now.

(300 * 1.304) + (1000 * 0.67) + (75 * 1.90)

= 391.20 + 670 + 142.50

= 1203.70

I have a total of $1203.70 now.


c

 The GBP/EUR rate is quoted here.

i) I will have to look for a higher bid rate because this the rate at which dealers are ready to buy the currency.

The bank is offering a bid rate of 0.665 which is higher than the other offer so I would select it.

ii) 300 * 0.665 = 199.50

I will receive EUR199.50